OpenStudy (anonymous):

A tuition scale for a state university shows an increase of $1,100 each year. If the initial cost was $27,000. What will the cost of tuition be in twenty two years? In complete sentences, explain how an arithmetic sequence was applied? Please help me to understand?

OpenStudy (dumbcow):

each year you are adding $1,100 tuition is $27,000, $28,100, $29,200, $30,300 ... it forms an arithmetic sequence An = A1 +d(n-1) A1 = 27,000 d = 1100 n=22 tuition in 22 years = 27,000 + 1100(22-1)

OpenStudy (dumbcow):

By the time you get to the 22nd year $1100 will have been added 21 times thats why in the formula you multiply 1100 by 21 then add to the initial tuition of 27000

OpenStudy (anonymous):

So you complete tuition in 22 years = 27,000 + 1100(22-1) and that is your final answer?

OpenStudy (dumbcow):

yes that will be what the tuition will be in 22 years. the reason it forms an arithmetic sequence is because each year the same amount(1100) is added

OpenStudy (anonymous):

Oh okay, so an arithmetic sequence is when a constant number is added in this case the 1100?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

thank you very much! I have one more I am going to ask but I am going to ask it in a separate question and try to solve it myself first with the equation

OpenStudy (dumbcow):

ok great, i will check your answer if you like

OpenStudy (anonymous):

yes! that would be awesome. i'll post it in a seperate question once i get done trying to solve it.